OpenStudy (anonymous):

find the equation of the ellipse with foci on the x-axis, major axis 9 units long, minor axis 4 units long, and center at the origin

OpenStudy (anonymous):

Do u know the formula for major axis length and minor axis length?

OpenStudy (anonymous):

No I don't @SandeepReddy

OpenStudy (anonymous):

what is the standard equation of ellipse?

OpenStudy (anonymous):

x^2/a^2 + y^2/b^2=1 @SandeepReddy

OpenStudy (anonymous):



OpenStudy (anonymous):

General equation of an ellipse with center at (x1,y1) is\[\frac{ (x-x1)^2 }{ a^2 }+\frac{ (y-y1)^2 }{ b^2 }=1\]

OpenStudy (anonymous):

Here he given center (x1,y1) = (0,0) so what u said was correct for this pe

OpenStudy (anonymous):

prblm

OpenStudy (anonymous):

thank so much! @sandeepreddy

OpenStudy (anonymous):

Its my pleasure @teebabyy